OpenStudy (anonymous):

how to integrate (1/2)[e^(1/2) dt ?

zepdrix (zepdrix):

\[\huge \int\limits \frac{1}{2}e^{1/2}dt\]Like this? No t in the function?

OpenStudy (anonymous):

yeah its like that :/ except the 1/2 is outside

zepdrix (zepdrix):

Inside outside, blah blah doesn't matter :) hehe

OpenStudy (anonymous):

haha just clarifying

zepdrix (zepdrix):

Ok they're just trying to confuse you with this one, it's actually very simple. The e^1/2 is just a fancy fancy looking constant. Nothing more. He's just a constant wearing a fancy tuxedo trying to confuse you and make you think you have to integrate it. He's just a constant. We can pull him outside of the integral along side the 1/2.\[\huge \frac{1}{2}e^{1/2}\int\limits\limits dt\]

zepdrix (zepdrix):

Confused? :o

OpenStudy (anonymous):

i got it now thanks haha

zepdrix (zepdrix):

cool c: